OpenStudy (anonymous):

angle a and angle b are complementary angles. If angle a is six more than three times angle b. What are the measures of the angles?

OpenStudy (anonymous):

b=43.5 a=136.5

OpenStudy (anonymous):

let a=6+3b complementary = total angles is 90 a+b=90 (6+3b)+b=90 6+3b+b=90 6+4b=90 4b=90-6 4b=84 b=21 90-21=69 angle a = 69 angle b = 21

OpenStudy (anonymous):

no no!!! a+b=180!

OpenStudy (anonymous):

complementary angles are 90 supplementary angles are 180 please check wikipedia... http://en.wikipedia.org/wiki/Complementary_angles

OpenStudy (anonymous):

oh yes i make a mistake sorry :)